# thanks: Corresponding author

### v1

@@ Abstract
The deep connection between black hole thermodynamics and spacetime geometry remains a central focus of general relativity. While recent studies have revealed a precise correspondence for null orbits, given by $K=-\lambda^{2}$ between the Gaussian curvature $K$ and the Lyapunov exponent $\lambda$, its validity for timelike orbits had remained unknown. Our work introduces the massive particle surface (MPS) framework and constructs a new geometric quantity $\mathcal{G}$. We demonstrate that $\mathcal{G}\propto-\lambda^{2}$ on unstable timelike orbits, thus establishing the geometry-dynamics correspondence for massive particles. Crucially, near the first-order phase transition of a black hole, $\mathcal{G}$ displays synchronized multivalued behavior with the Lyapunov exponent $\lambda$ and yields a critical exponent $\delta=1/2$. Our results demonstrate that spacetime geometry encodes thermodynamic information, opening a new pathway for studying black hole phase transitions from a geometric perspective.
Introduction. Black holes, remarkable consequences of general relativity, exhibit rich thermodynamic phenomena Hawking (1975); Bekenstein (1973), including first-order phase transitions analogous to van der Waals fluids Kubiznak and Mann (2012). This finding challenges the conventional boundaries between gravity, thermodynamics, and statistical physics, raising a central question: How are the thermodynamic properties of black holes fundamentally encoded in the geometric structure of spacetime?
@@
This work answers both questions by introducing the massive particle surface (MPS) framework. We construct a new geometric quantity $\mathcal{G}$ to demonstrate that, for neutral massive particles on unstable timelike orbits, the relation $\mathcal{G}\propto-\lambda^{2}$ holds. This result successfully extends the geometry-dynamics correspondence from massless to massive particles. Crucially, in the spinodal region of the black hole first-order phase transition, both $\mathcal{G}$ and the Lyapunov exponent $\lambda$ simultaneously exhibit multivalued behavior, with a critical exponent of $\delta=1/2$ consistent with mean-field theory. Our work establishes spacetime geometry as a unified foundation connecting dynamic chaos and thermodynamic phase transitions in black holes, thereby enabling a geometric perspective for probing their thermodynamics. We set $G=c=k_{B}=\hbar=1$ in this paper.
Timelike Unstable Orbits and Lyapunov Exponent. The Lyapunov exponent $\lambda$ quantifies the chaotic behavior of particle motion on timelike circular orbits near a black hole. First, we briefly review the derivation of this exponent for unstable timelike orbits.
Consider a $(3+1)$-dimensional spherically symmetric metric
@@
where $d\Omega^{2}$ is the unit 2-sphere, and f,gf,~g are $C^{P}$ smooth functions of $r$ with $P\geq 2$. For timelike orbits, the Lagrangian is
$\mathcal{L}=g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=1.$ (2)
The two conserved quantities of the particle, its energy $\mathcal{E}$ and angular momentum $L$, are given by
@@
We expand the Lagrangian around the unstable circular orbit located at $r=r_{0}$, which leads to
@@
and define the Lyapunov exponent as
$\lambda=\sqrt{\frac{1}{2(\dot{t})^{2}}V_{\text{eff}}^{\prime\prime}}\bigg|_{r=r_{0}},$ (5)
where $r_{0}$ is determined by the conditions $V_{\text{eff}}(r_{0})=V_{\text{eff}}^{\prime}(r_{0})=0$. When $f=g$, we obtain
@@
Jacobi Metric and the Limitation of Gaussian Curvature. For a $(3+1)$-dimensional spherically symmetric metric as in Eq. (1), a precise correspondence between orbital dynamics and geometry for unstable null orbits Gallo and Mädler (2025)
@@
MPS Geometry. We briefly review the fundamental concepts and properties of the MPS defined in Refs. Kobialko et al. (2022); Bogush et al. (2023, 2024), and establish the connection between the associated geometric quantity $\mathcal{G}$ and the Lyapunov exponent.
@@
On the MPS, consider an orthonormal basis constructed from the timelike Killing vector $\tilde{k}^{\mu}$ on $\Sigma$ and a set of d2d-2 linearly independent vectors $\tau^{\alpha}_{(i)}$ defined on $\Sigma$, satisfying
@@
Now consider a general $(3+1)$-dimensional spherically symmetric metric
$ds^{2}=-fdt^{2}+\frac{1}{g}dr^{2}+hd\Omega^{2},$ (28)
where $f$, $g$, and $h$ are $C^{P}$ smooth functions of $r$ with $P\geq 2$. For this metric, we obtain
@@
Solving Eq. (31) under this condition yields the radius $r_{LR}$ of the null unstable circular orbit, which is fully consistent with the result obtained via the dynamical approach. This confirms that in the massless limit, the MPS naturally reduces to the photon surface. For $h=r^{2}$, one can find Bermúdez-Cárdenas and Andino (2025)
$\frac{d}{dr}\left(\frac{\tilde{k}^{2}\chi_{\tau}}{W}\right)=\frac{d}{dr}\left(\frac{\mathcal{E}^{2}}{m^{2}}\right)=\frac{3\frac{ff^{\prime}}{r}-2(f^{\prime})^{2}+ff^{\prime\prime}}{r(2f-rf^{\prime})^{2}}.$ (32)
@@
$\mathcal{G}=-\frac{2}{r(2f-rf^{\prime})^{2}}\lambda^{2},$ (33)
@@
Results. Eq. (33) is now employed to demonstrate the multivalued behavior of $\mathcal{G}$ near black hole first-order phase transition points, thereby showing its effectiveness in probing such phase transitions in the timelike case. We consider the Hayward-Letelier-AdS black hole as a specific example.
We begin by reviewing relations in black hole thermodynamics. For a $(3+1)$-dimensional spherically symmetric black hole described by the metric in Eq. (28), when $f=g$ and $h=r^{2}$, the Hawking temperature is given by
@@
where $M$ is the ADM mass of the black hole and $S=\pi r_{+}^{2}$ is the entropy. When a black hole undergoes a first-order phase transition, the $F(T)$ curve exhibits a swallowtail structure, corresponding to the existence of three black hole solutions: large, intermediate, and small black holes. The critical point is determined by the conditions
@@
Figure 1: Phase transition signatures in dynamics and geometry for timelike orbits, with $\tilde{g}_{c}=0.09002$, $a=0.6$, $L=20\ell$. Left column ($\tilde{g}=0.061538$, $\tilde{g}<\tilde{g}_{c}$): (a) free energy $\tilde{F}_{HL}$ versus $\tilde{T}_{HL}$, (c) Lyapunov exponent $\lambda_{HL}$ versus $\tilde{T}_{HL}$, (e) geometric quantity $|\mathcal{G}_{HL}|$ versus $\tilde{T}_{HL}$ (log scale). Right column ($\tilde{g}=0.092308$, $\tilde{g}>\tilde{g}_{c}$): (b) $\tilde{F}_{HL}$ versus $\tilde{T}_{HL}$, (d) $\lambda_{HL}$ versus $\tilde{T}_{HL}$, (f) $|\mathcal{G}_{HL}|$ versus $\tilde{T}_{HL}$ (log scale). The synchronized multivalued behavior of $\lambda_{HL}$ and $|\mathcal{G}_{HL}|$ in the spinodal region $\tilde{T}_{HL}\in(\tilde{T}_{1},\,\tilde{T}_{2})$ corresponds to the swallowtail structure in the free energy, with the phase transition occurring at $\tilde{T}_{p}$. All quantities become monotonic in the absence of a phase transition.
@@
Critical Exponents. Expanding $\mathcal{G}$ near the critical point $\tilde{r}_{c}$ using the standard method Banerjee and Roychowdhury (2012a, b) reveals a scaling behavior consistent with the mean-field prediction
$\Delta\mathcal{G}=(\mathcal{G}_{+}-\mathcal{G}_{c})\sim\left(\tilde{T}-\tilde{T}_{c}\right)^{\delta},$ (40)
[…]
We begin by expanding $\mathcal{G}$ near the critical point $\tilde{r}_{c}$
$\mathcal{G}_{+}=\mathcal{G}_{c}+\left(\frac{\partial\mathcal{G}}{\partial\tilde{r}_{+}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{2}.$ (41)
Introducing the small parameter $\xi\ll 1$ through
$\tilde{r}_{+}=\tilde{r}_{c}(1+\xi),$ (42)
we expand the Hawking temperature
$\tilde{T}=\tilde{T}_{c}+\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})^{2}+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{3},$ (43)
which leads to
$\tilde{r}_{+}-\tilde{r}_{c}\approx(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}.$ (44)
Substituting Eq. (44) into Eq. (41) leads to
$\mathcal{G}_{+}-\mathcal{G}_{c}=\left(\frac{\partial\mathcal{G}}{\partial\tilde{r}_{+}}\right)_{c}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}},$ (45)
which demonstrates the scaling relation Eq. (40), with the critical exponent $\delta=1/2$. This result is consistent with previous studies of both spherically symmetric and non-spherically symmetric black holes Guo et al. (2022); Lyu et al. (2024); Shukla et al. (2024); Chen et al. (2025); Yang et al. (2025).
@@
Conclusion. In this work, we have systematically extended the correspondence between geometry and dynamics from unstable null orbits to the timelike case and uncovered their fundamental distinction. Within the framework of the MPS theory, we constructed a new geometric quantity $\mathcal{G}$ and demonstrated that $\mathcal{G}\propto-\lambda^{2}$ holds for unstable timelike orbits, where $\lambda$ is the Lyapunov exponent characterizing the orbital chaos. This finding fills a crucial theoretical gap by establishing this correspondence specifically for massive particles.
@@
Moreover, we have found that the geometric quantity $\mathcal{G}$ exhibits multivalued behavior near the black hole first-order phase transition, synchronizing with the Lyapunov exponent $\lambda$. The extracted critical exponent $\delta=1/2$ agrees with the mean-field theory. This reveals that spacetime geometry itself encodes thermodynamic phase transition information, thereby establishing a geometric framework for probing black hole thermodynamics in the timelike case.

### v2

@@ Abstract
The deep connection between black hole thermodynamics and spacetime geometry remains a central focus of general relativity. While recent studies have revealed a precise correspondence for null orbits, given by $K=-\lambda^{2}$ between the Gaussian curvature $K$ and the Lyapunov exponent $\lambda$, its validity for timelike orbits had remained unknown. Our work introduces the massive particle surface (MPS) framework and constructs a new geometric quantity $\mathcal{G}$. We demonstrate that $\mathcal{G}\propto-\lambda^{2}$ on unstable timelike orbits, thus establishing the geometry-dynamics correspondence for massive particles. Crucially, near the first-order phase transition of a black hole, $\mathcal{G}$ displays synchronized multivalued behavior with the Lyapunov exponent $\lambda$ and yields a critical exponent $\delta=1.0244$. Our results demonstrate that spacetime geometry encodes thermodynamic information, opening a new pathway for studying black hole phase transitions from a geometric perspective.
I Introduction
Black holes, remarkable consequences of general relativity, exhibit rich thermodynamic phenomena Hawking (1975); Bekenstein (1973), including first-order phase transitions analogous to van der Waals fluids Kubiznak and Mann (2012). This finding challenges the conventional boundaries between gravity, thermodynamics, and statistical physics, raising a central question: How are the thermodynamic properties of black holes fundamentally encoded in the geometric structure of spacetime?
@@ I Introduction
This work answers both questions by introducing the massive particle surface (MPS) framework. We construct a new geometric quantity $\mathcal{G}$ to demonstrate that, for neutral massive particles on unstable timelike orbits, the relation $\mathcal{G}\propto-\lambda^{2}$ holds. This result successfully extends the geometry-dynamics correspondence from massless to massive particles. Crucially, in the spinodal region of the black hole first-order phase transition, both $\mathcal{G}$ and the Lyapunov exponent $\lambda$ simultaneously exhibit multivalued behavior. By examining the critical behavior of these quantities near the phase transition point, we find that their critical exponents deviate from the value $1/2$ observed in the Reissner–Nordström black hole. This indicates that regular black holes exhibit richer critical behavior in their geometry and dynamics than their singular counterparts. Our work establishes spacetime geometry as a unified foundation connecting dynamic chaos and thermodynamic phase transitions in black holes, thereby enabling a geometric perspective for probing their thermodynamics. We set $G=c=k_{B}=\hbar=1$ in this paper.
II Timelike Unstable Orbits and Lyapunov Exponent
The Lyapunov exponent $\lambda$ quantifies the chaotic behavior of particle motion on timelike circular orbits near a black hole. First, we briefly review the derivation of this exponent for unstable timelike orbits.
Consider a 4-dimensional static spherically symmetric metric
@@ II Timelike Unstable Orbits and Lyapunov Exponent
where $d\Omega^{2}=r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ is the unit 2-sphere, and f,gf,~g are $C^{P}$ smooth functions of $r$ with $P\geq 2$. For timelike orbits, the Lagrangian is
$\mathcal{L}=g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=-1,$ (2)
where the dot denotes a derivative with respect to affine parameter. The two conserved quantities of the particle, its energy $\mathcal{E}$ and angular momentum $L$, are given by
@@
We expand the Lagrangian around the unstable circular orbit located at $r=r_{0}$, and restrict our attention to the metric in the equatorial plane ($\theta=\frac{\pi}{2}$), which leads to
@@
the prime notation denotes differentiation with respect to the radial coordinate $r$, and define the Lyapunov exponent as
$\lambda=\sqrt{\frac{1}{2(\dot{t})^{2}}V_{\text{eff}}^{\prime\prime}}\bigg|_{r=r_{0}}.$ (5)
With
$V_{\text{eff}}=g\left(\frac{E^{2}}{f}-\frac{L^{2}}{r^{2}}-1\right),$ (6)
where $r_{0}$ is determined by the conditions $V_{\text{eff}}(r_{0})=V_{\text{eff}}^{\prime}(r_{0})=0$, and $\varepsilon=r-r_{0}$. When $f=g$, we obtain
@@
III Jacobi Metric and the Limitation of Gaussian Curvature
For a $4$-dimensional spherically symmetric metric as in Eq. (1), a precise correspondence between orbital dynamics and geometry for unstable null orbits Gallo and Mädler (2025)
@@ III Jacobi Metric and the Limitation of Gaussian Curvature
IV MPS Geometry
We briefly review the fundamental concepts and properties of the MPS defined in Refs. Kobialko et al. (2022); Bogush et al. (2023, 2024), and establish the connection between the associated geometric quantity $\mathcal{G}$ and the Lyapunov exponent.
@@ IV MPS Geometry
This is the special uncharged case of Ref. Kobialko et al. (2022). On the MPS, consider an orthonormal basis constructed from the timelike Killing vector $\tilde{k}^{\mu}$ on $\Sigma$ and a set of d2d-2 linearly independent vectors $\tau^{\alpha}_{(i)}$ defined on $\Sigma$, satisfying
@@
Now consider a 4-dimensional static spherically symmetric metric as in Eq. (1). For this metric, we obtain
@@
Solving Eq. (31) under this condition yields the radius $r_{LR}$ of the null unstable circular orbit, which is fully consistent with the result obtained via the dynamical approach. This confirms that in the massless limit, the MPS naturally reduces to the photon surface. For $h=r^{2}$, one can find
$\frac{d}{dr}\left(\frac{\tilde{k}^{2}\chi_{\tau}}{W}\right)=\frac{d}{dr}\left(\frac{\mathcal{E}^{2}}{m^{2}}\right)=\frac{2fr}{(2f-rf^{\prime})^{2}}\left(3\frac{ff^{\prime}}{r}-2f^{\prime 2}+ff^{\prime\prime}\right).$ (32)
@@
$\mathcal{G}=-\frac{4fr}{(2f-rf^{\prime})^{2}}\lambda^{2},$ (33)
@@
V Results
Eq. (33) is now employed to demonstrate the multivalued behavior of $\mathcal{G}$ near black hole first-order phase transition points, thereby showing its effectiveness in probing such phase transitions in the timelike case. We consider the Hayward-Letelier-AdS black hole as a specific example.
We begin by reviewing relations in black hole thermodynamics. For a $4$-dimensional spherically symmetric black hole described by the metric in Eq. (1), when $f=g$ and $h=r^{2}$, the Hawking temperature is given by
@@ V Results
where $M$ is the ADM mass of the black hole and $S=\pi r_{+}^{2}$ is the entropy. When a black hole undergoes a first-order phase transition, the $F(T)$ curve exhibits a swallowtail structure, corresponding to the existence of three black hole solutions: large, intermediate, and small black holes. The critical point is an inflection point determined by the conditions
@@
(a) $\tilde{F}_{HL}-\tilde{T}_{HL}$
(b) $\tilde{F}_{HL}-\tilde{T}_{HL}$
(c) $\lambda_{HL}-\tilde{T}_{HL}$ (Null)
(d) $\lambda_{HL}-\tilde{T}_{HL}$
(e) $|\mathcal{G}_{HL}|-\tilde{T}_{HL}$
(f) $|\mathcal{G}_{HL}|-\tilde{T}_{HL}$
Figure 1: Phase transition signatures in dynamics and geometry for timelike orbits, with $\tilde{g}_{c}=0.1042$, $a=0.6$, $L=20\ell$. Left column ($\tilde{g}=0.0615$, $\tilde{g}<\tilde{g}_{c}$): (a) free energy $\tilde{F}_{HL}$ versus $\tilde{T}_{HL}$, (c) Lyapunov exponent $\lambda_{HL}$ versus $\tilde{T}_{HL}$, (e) geometric quantity $|\mathcal{G}_{HL}|$ versus $\tilde{T}_{HL}$ (log scale). Right column ($\tilde{g}=0.1276$, $\tilde{g}>\tilde{g}_{c}$): (b) $\tilde{F}_{HL}$ versus $\tilde{T}_{HL}$, (d) $\lambda_{HL}$ versus $\tilde{T}_{HL}$, (f) $|\mathcal{G}_{HL}|$ versus $\tilde{T}_{HL}$ (log scale). The synchronized multivalued behavior of $\lambda_{HL}$ and $|\mathcal{G}_{HL}|$ in the spinodal region $\tilde{T}_{HL}\in(\tilde{T}_{1},\,\tilde{T}_{2})$ corresponds to the swallowtail structure in the free energy, with the phase transition occurring at $\tilde{T}_{p}$. All quantities become monotonic in the absence of a phase transition.
@@
VI Critical Exponents
Critical exponents are used to describe the behavior of physical quantities of thermodynamic systems near the critical point and serve as important physical quantities for characterizing phase transitions. At the phase transition point, the small and large black hole branches coexist, leading to discontinuities in both the timelike Lyapunov exponent $\lambda$ and the geometric quantity $\mathcal{G}$. This section examines the continuity of $\lambda$ and $\mathcal{G}$ at the phase transition point $T_{p}$ for the Hayward–Letelier–AdS black hole, as well as their discontinuity near $T_{p}$. We investigate whether they can serve as order parameters and check whether their critical behavior aligns with the predictions of mean-field theory.
We define the timelike Lyapunov exponents for the small and large black hole branches within the spinodal region as $\lambda_{S}$ and $\lambda_{L}$, respectively, and the corresponding geometric quantities as $\mathcal{G}_{S}$ and $\mathcal{G}_{L}$. Their differences are defined as $\Delta\lambda$ and $\Delta\mathcal{G}$. When the critical point $T_{p}=T_{c}$, we have $\Delta\lambda=\Delta\mathcal{G}=0$. We will fit the numerical coefficients of $\Delta\lambda$ and $\Delta\mathcal{G}$ according to the following form Wei and Liu (2018)
$\frac{\Delta\lambda}{\lambda_{c}},\frac{\Delta\mathcal{G}}{\mathcal{G}_{c}}\sim\alpha(t-1)^{\beta}.$ (40)
Here, $t=T_{p}/T_{c}$, $\alpha$ and $\beta$ are real numbers, the subscript “cc” denotes the value of $\lambda$ and $\mathcal{G}$ at the critical point.
Figure 2: The reduced timelike Lyapunov exponents ${\Delta\lambda}/{\lambda_{c}}$ and geometric quantity ${\Delta\mathcal{G}}/{\mathcal{G}_{c}}$ for the coexistence small and large black holes. (a) ${\Delta\lambda}/{\lambda_{c}}$ and (b) ${\Delta\mathcal{G}}/{\mathcal{G}_{c}}$ versus reduced temperature $t$ for a Hayward-Letelier-AdS black hole.
As shown in Fig. 2, we plot the normalized differences ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta\mathcal{G}}/{\mathcal{G}_{c}}$ as functions of the reduced temperature $t$. When $T_{p}=T_{c}$, ${\Delta\lambda}/{\lambda_{c}}={\Delta\mathcal{G}}/{\mathcal{G}_{c}}=0$. When $T_{p}$ deviates from $T_{c}$, both ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta\mathcal{G}}/{\mathcal{G}_{c}}$ take non-zero values, indicating the emergence of discontinuity. Finally, by fitting the curves of $\Delta\lambda/\lambda_{c}(t)$ and $\Delta\mathcal{G}/\mathcal{G}_{c}(t)$, we obtain the following scaling behavior
$\frac{\Delta\lambda}{\lambda_{c}}\simeq 10^{0.8286}(t-1)^{0.5918},$ (41)
and
$\frac{\Delta\mathcal{G}}{\mathcal{G}_{c}}\simeq 10^{14.6054}(t-1)^{1.0244}.$ (42)
These results indicate that the critical exponents of ${\Delta\lambda}/{\lambda_{c}}$ ($0.5918$) and ${\Delta\mathcal{G}}/{\mathcal{G}_{c}}$ ($1.0244$) deviate from the value of $1/2$ observed in the Reissner-Nordström black hole, suggesting that regular black holes exhibit richer critical behavior in their geometry and dynamics than their singular counterparts. Given Eq. (33), the geometric quantity $\mathcal{G}$ is proportional to $\lambda^{2}$, but the proportionality coefficient involves metric functions. Therefore, the critical exponent of ${\Delta\mathcal{G}}/{\mathcal{G}_{c}}$ is not simply twice that of ${\Delta\lambda}/{\lambda_{c}}$ ($1.1836$), but is instead influenced by the scaling of the coefficient itself. This leads to a deviation of the critical exponent of ${\Delta\mathcal{G}}/{\mathcal{G}_{c}}$ from the expected value $1.1836$. This result further reveals the complex coupling between geometric and dynamic quantities in regular black holes.
VII Conclusion
@@
In this work, we have systematically extended the correspondence between geometry and dynamics from unstable null orbits to the timelike case and uncovered their fundamental distinction. Within the framework of the MPS theory, we constructed a new geometric quantity $\mathcal{G}$ and demonstrated that $\mathcal{G}\propto-\lambda^{2}$ holds for unstable timelike orbits, where $\lambda$ is the Lyapunov exponent characterizing the orbital chaos. This finding fills a crucial theoretical gap by establishing this correspondence specifically for massive particles.
@@
Moreover, we have found that the geometric quantity $\mathcal{G}$ exhibits multivalued behavior near the black hole first-order phase transition, synchronizing with the Lyapunov exponent $\lambda$. This reveals that spacetime geometry itself encodes thermodynamic phase transition information, thereby establishing a geometric framework for probing black hole thermodynamics in the timelike case. Furthermore, we find that the critical behaviors of both $\mathcal{G}$ and $\lambda$ deviate from the mean-field expectation, indicating that the geometric and dynamic properties of regular black holes differ from those of their singular counterparts. This result also provides a basis for further investigation into this distinction.